# On the optimality of MCMC integration for functions with unbounded second moment

### v1

@@ 3 Proofs
Set $T=\operatorname{\text{Id}}-K$. Note that by assumption $\hat{h}\in L^{2}_{0}(\pi)$ and
$h=TT^{-1}h=T\hat{h}=\hat{h}-K\hat{h}.$
Thus $\mathbb{P}_{\nu}$-almost surely
$S_{n}h$ $=\frac{1}{n}\sum_{j=1}^{n}\hat{h}(X_{j})-K\hat{h}(X_{j})$
$=\frac{1}{n}\sum_{j=1}^{n}\left(\hat{h}(X_{j+1})-K\hat{h}(X_{j})\right)+\frac{1}{n}\left(\hat{h}(X_{1})-\hat{h}(X_{n+1})\right)$
$=\frac{1}{n}M_{n}+\frac{1}{n}\left(\hat{h}(X_{1})-\hat{h}(X_{n+1})\right).$
Since $(X_{n})_{n\in\mathbb{N}}$ is a Markov chain with kernel $K$ for any $n\geq 2$ holds
$\mathbb{E}_{\nu}[M_{n}|\mathcal{F}_{n-1}]=M_{n-1}+\mathbb{E}_{\nu}\left[\hat{h}(X_{n+1})-K\hat{h}(X_{n})\;\big{|}\;\mathcal{F}_{n-1}\right]=M_{n-1}.$
It remains to show the bound for the differences $\Delta_{j}$, which also implies that $\mathbb{E}_{\nu}|M_{n}|<\infty$ for any $n\in\mathbb{N}$ and hence that $(M_{n})_{n\in\mathbb{N}}$ is a martingale. Thereto observe that $\|\hat{h}\|_{L^{2}(\pi)}\leq s\left\|h\right\|_{L^{2}(\pi)}$ and that $\left\|K\right\|_{L^{1}(\pi)\to L^{1}(\pi)}=1$. Hence, Hoelder’s inequality shows that for any $j\in\mathbb{N}$ holds
$\mathbb{E}_{\nu}\left[|\hat{h}(X_{j+1})|^{2}\right]=\int_{\mathcal{X}}K^{j}|\hat{h}|^{2}(x)\frac{{\rm d}\nu}{{\rm d}\pi}(x)\pi({\rm d}x)\leq\left\|K^{j}|\hat{h}|^{2}\right\|_{L^{1}(\pi)}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}$
$\leq\|\hat{h}\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}\leq s^{2}\|h\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}.$
Similarly, $\mathbb{E}_{\nu}[(K\hat{h}(X_{j}))^{2}]\leq\mathbb{E}_{\nu}[K\hat{h}^{2}(X_{j})]\leq s^{2}\|h\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}$. The bound for $\mathbb{E}_{\nu}|\Delta_{j}|^{2}$ now follows by using the inequality $(\alpha+\beta)^{2}\leq 2(\alpha^{2}+\beta^{2})$, being valid for any $\alpha,\beta\in\mathbb{R}$. ∎
Based on the above martingale decomposition of Lemma 3.1 we now show a bound for the mean squared error.
Proposition 3.2.
Let Assumption 2.1 be true and assume that $\nu$ is absolutely continuous w.r.t. $\pi$ with Radon-Nikodym derivative $\frac{{\rm d}\nu}{{\rm d}\pi}\in L^{\infty}(\pi)$. Then, for any $f\in L^{2}(\pi)$ we have that
$\mathbb{E}_{\nu}\left[\left|S_{n}f-\pi(f)\right|^{2}\right]\leq 6s^{2}\left(1+\frac{1}{n}\right)\frac{\left\|f\right\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}}{n}.$
Proof.
Let $h=f-\pi(f)\in L^{2}_{0}(\pi)$. Lemma 3.1 and the inequality $(\alpha+\beta+\gamma)^{2}\leq 3(\alpha^{2}+\beta^{2}+\gamma^{2})$, which is true for any real numbers $\alpha,\beta,\gamma$, yield
$\mathbb{E}_{\nu}\left[|S_{n}f-\pi(f)|^{2}\right]=\mathbb{E}_{\nu}\left[(S_{n}h)^{2}\right]\leq\frac{3}{n^{2}}\mathbb{E}_{\nu}\left[M_{n}^{2}\right]+\frac{3}{n^{2}}\mathbb{E}_{\nu}[\hat{h}(X_{1})^{2}+\hat{h}(X_{n+1})^{2}],$
with $M_{n},\hat{h}$ as specified in Lemma 3.1. By the martingale property (and again using the notation/results of Lemma 3.1) it follows that
$\mathbb{E}_{\nu}\left[M_{n}^{2}\right]=\mathbb{E}_{\nu}\left[\sum_{j=1}^{n}\Delta_{j}^{2}\right]\leq\sum_{j=1}^{n}\mathbb{E}_{\nu}\left[\Delta_{j}^{2}\right]\leq 2ns^{2}\|h\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}.$
Furthermore, the same arguments as in the proof of Lemma 3.1 reveal
$\mathbb{E}_{\nu}[\hat{h}(X_{1})^{2}+\hat{h}(X_{n+1})^{2}]\leq 2s^{2}\|h\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}.$
Combining these estimates with $\left\|h\right\|_{L^{2}(\pi)}=\left\|f\right\|_{L^{2}(\pi)}^{2}-\pi(f)^{2}\leq\left\|f\right\|_{L^{2}(\pi)}^{2}$ yield the desired result. ∎
Remark 3.3.
Note that Proposition 3.2 can be seen as extension of two results in [Rud12]:
  • It extends Theorem 3.34 to the case where $K$ is not required to be reversible [^2] Note that $L^{1}$-exponential convergence implies $L^{2}$-exponential convergence, e.g. by an application of the Riesz-Thorin interpolation theorem, see [Gra14, Theorem 1.3.4]. Thus Assumption 2.1 holds true in this case..
  • It extends Theorem 3.41 to the $L^{2}$ setting.
Finally, we turn to the proof of our main result.
@@
We write $L^{q}(\mathbb{P}_{\nu})$ for the set of random variables $X$ with finite $\left\|X\right\|_{L^{q}(\mathbb{P}_{\nu})}^{q}=\mathbb{E}_{\nu}|X|^{q}$. For $n\in\mathbb{N}$ consider the linear mapping $T_{n}f=S_{n}f-\pi(f)$ and note that
$\mathbb{E}_{\nu}|T_{n}f|\leq\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\nu}|f(X_{j})|+\pi(|f|)=\frac{1}{n}\sum_{j=1}^{n}\int_{\mathcal{X}}K^{j-1}|f|(x)\frac{{\rm d}\nu}{{\rm d}\pi}(x)\pi({\rm d}x)+\pi(|f|).$
By Hoelder’s inequality and the fact that $\left\|K\right\|_{L^{1}(\pi)\to L^{1}(\pi)}=1$ it follows that $\left\|T_{n}\right\|_{L^{1}(\pi)\to L^{1}(\mathbb{P}_{\nu})}\leq 1+\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}$.
Additionally, by Proposition 3.2 holds
$\left\|T_{n}\right\|_{L^{2}(\pi)\to L^{2}(\mathbb{P}_{\nu})}\leq\frac{1}{\sqrt{n}}\left(12s^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}\right)^{1/2}.$
An application of the Riesz-Thorin interpolation theorem, see [Gra14, Theorem 1.3.4] in the setting $p_{0}=q_{0}=1$; $p_{1}=q_{1}=2$ and $\theta=2-2/p$, yields that
$\left\|T_{n}\right\|_{L^{p}(\pi)\to L^{p}(\mathbb{P}_{\nu})}\leq\left\|T_{n}\right\|_{L^{1}(\pi)\to L^{1}(\mathbb{P}_{\nu})}^{1-\theta}\left\|T_{n}\right\|_{L^{2}(\pi)\to L^{2}(\mathbb{P}_{\nu})}^{\theta},$
which, together with the above estimates, implies the result. ∎
@@ Acknowledgements
The author would like to thank Mareike Hasenpflug and Daniel Rudolf for valuable discussions which helped to significantly improve the presentation of the paper. JH is supported by the DFG within project 432680300 – SFB 1456 subproject B02.

### v2

@@ 3 Proofs
Since $\mathbb{P}_{\nu}$ and $\mathbb{P}_{\pi}$ only differ by the choice of the initial distribution of $(X_{n})_{n\in\mathbb{N}}$ we have
$\mathbb{E}_{\nu}[\left|S_{n}f-\pi(f)\right|^{p}]=\mathbb{E}_{\pi}\left[\frac{{\rm d}\nu}{{\rm d}\pi}(X_{1})\left|S_{n}f-\pi(f)\right|^{p}\right]\leq\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}\mathbb{E}_{\pi}[\left|S_{n}f-\pi(f)\right|^{p}].$
Hence, the result follows from Proposition 3.2. ∎
Finally, we turn to the proof of Theorem 2.3
@@
By the same change of measure argument as before and Hoelder’s inequality
$\mathbb{E}_{\nu}[\left|S_{n}f-\pi(f)\right|]=\mathbb{E}_{\pi}\left[\frac{{\rm d}\nu}{{\rm d}\pi}(X_{1})\left|S_{n}f-\pi(f)\right|\right]\leq\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{q}(\pi)}\mathbb{E}_{\pi}[\left|S_{n}f-\pi(f)\right|^{p}]^{1/p}.$
Now the desired result is a consequence of Proposition 3.2. ∎
@@ Acknowledgements
The author would like to thank Mareike Hasenpflug and Daniel Rudolf for valuable discussions which helped to significantly improve the presentation of the paper. Moreover, the author thanks two anonymous referees for reading the paper carefully, and suggesting a simplification for Section 3. JH is supported by the DFG within project 432680300 – SFB 1456 subproject B02.